# EO-1: Interleaved Vision-Text-Action Pretraining for General Robot Control

### v4

@@ 4.3 Benchmark Evaluation Results
RT-1 (Brohan et al., 2022) 85.7% 44.2% 73.0% 6.5% 52.4%
OpenVLA (Kim et al., 2024) 16.3% 46.2% 35.6% 0% 24.5%
TraceVLA (Zheng et al., 2024) 28.0% 53.7% 57.0% 0% 34.7%
SpatialVLA (Qu et al., 2025) 86.0% 77.9% 57.4% 0% 55.3%
Magma (Yang et al., 2025) 75.0% 53.0% 58.9% 8.3% 48.8%
$\pi_{0}$-FAST (Pertsch et al., 2025) 75.3% 67.5% 42.9% 0 46.4%
RT-1-X (OpenX-Embodiment et al., 2024) 56.7% 31.7% 59.7% 40.7% 47.2%
RT-2-X (OpenX-Embodiment et al., 2024) 78.7% 77.9% 25.0% 7.4% 47.3%
Octo-Base (Octo et al., 2024) 17.0% 4.2% 22.7% 0 11.0%
RoboVLM Li et al. (2024a) 77.3% 61.7% 43.5% 24.1% 51.7%
$\pi_{0}$ (Black et al., 2024) 97.9% 78.7% 62.25% 46.6% 71.4%
ThinkAct (Huang et al., 2025) 92.0% 72.4% 50.0%
MolmoAct (Lee et al., 2025) 77.7% 77.1% 60.0%
@@
RT-1 (Brohan et al., 2022) 89.8% 50.0% 32.3% 2.6% 43.7%
OpenVLA (Kim et al., 2024) 54.5% 47.7% 17.7% 0.0% 30.0%
TraceVLA (Zheng et al., 2024) 60.0% 56.4% 31.0% 0% 36.9%
SpatialVLA (Qu et al., 2025) 88.0% 72.7% 41.8% 6.3% 52.2%
Magma (Yang et al., 2025) 68.6% 78.5% 59.0% 24.0% 57.5%
$\pi_{0}$-FAST (Pertsch et al., 2025) 77.6% 68.2% 31.3% 0 44.3%
RT-1-X (OpenX-Embodiment et al., 2024) 49.0% 32.3% 29.4% 10.1% 30.2%
RT-2-X (OpenX-Embodiment et al., 2024) 82.3% 79.2% 35.3% 20.6% 54.4%
Octo-Base (Octo et al., 2024) 0.6% 3.1% 1.1% 0 1.2%
RoboVLM Li et al. (2024a) 75.6% 60.0% 10.6% 0 36.6%
$\pi_{0}$ (Black et al., 2024) 90.1% 80.7% 27.6% 20.5% 54.7%
ThinkAct (Huang et al., 2025) 84.0% 63.8% 47.6%
@@
Table 4: Performance comparison with state-of-the-art policies on SimplerEnv Benchmark (Li et al., 2024b).
Robot Control. We evaluated manipulation performance on two simulation benchmarks: LIBERO (Liu et al., 2023a)and SimplerEnv (Li et al., 2024b). LIBERO spans four subsets, Spatial, Object, Goal, and Long, designed to test generalization across spatial layouts, object categories, goal semantics, and long-horizon planning. SimplerEnv includes Google-VM (Visual Matching), Google-VA (Visual Aggregation), and WindoWX setups with controlled variations in color, texture, lighting, and camera pose, assessing robustness under visual distribution shift. We compare against leading robot foundation models, including the RT series (Brohan et al., 2022, OpenX-Embodiment et al., 2024), Octo (Octo et al., 2024), OpenVLA (Kim et al., 2024), the $\pi$-series (Black et al., 2024, Pertsch et al., 2025), and ThinkAct (Huang et al., 2025), among others.
As shown in Table˜3 and Table˜4, our method achieves superior performance across various robotic manipulation tasks. On the LIBERO benchmark, it attains an average success rate of 98.2%, outperforming recent state-of-the-art models, including OpenVLA-OFT (Kim et al., 2025), $\pi_{0}$ (Black et al., 2024), and GR00t-N1 (Bjorck et al., 2025) by 1.1%, 4.0%, and 4.3%, respectively, demonstrating robust generalization across spatial, semantic, and long-horizon tasks. On the SimplerEnv benchmarks, it surpasses $\pi_{0}$ (Black et al., 2024) by 3.5%, 5.1%, and 8.3% in the WidowX, Google-VM, and Google-VA variants, respectively, achieving the highest success rates of 72.7%, 76.5%, and 63.0%. These results are obtained with lightweight fine-tuning on a modest mixed-modality dataset, underscoring the method’s data efficiency, dexterous control, and precise visuomotor grounding.
@@ 4.5 Mastering Diverse Manipulations on Multiple Embodiments
In this section, we demonstrate that EO -1 can perform a wide range of dexterous manipulation tasks across various robotic platforms, showcasing its robustness and adaptability. We evaluate its performance on both short-horizon and long-horizon tasks using different robots, including Franka Panda, WidowX 250 S, Agibot G-1, and Lerobot SO100. These tasks range from basic pick-and-place to complex articulated object manipulation and long-horizon tasks. We compare EO -1 against state-of-the-art baselines, including $\pi_{0}$-Fast, $\pi_{0}$, and GR00T-N1.5 (Pertsch et al., 2025, Black et al., 2024, Bjorck et al., 2025), highlighting its superior performance across these varied tasks.
@@ 4.7 Emerging Open-world Embodied Generalization
Against strong baselines GR00T-N1.5 and $\pi_{0}$ (Pertsch et al., 2025, Black et al., 2024, Bjorck et al., 2025), EO -1 achieves an average completion score of 87.0%, outperforming GR00T-N1.5 (75.0%) and $\pi_{0}$ (66.0%). Performance gains are consistent across embodiments and task types, with EO -1 maintaining high accuracy even under precise spatial references and unseen object–instruction combinations, demonstrating robustness to both language and visual distribution shifts. We observe that while $\pi_{0}$ executes tasks quickly and accurately, it struggles to follow instructions, often prioritizing objects closer to itself instead of adhering to the specified commands. This behavior may be due to the negative impact of pre-training on the entire backbone, which can lead to suboptimal decision-making in dynamic environments. In contrast, GR00T-N1.5, which freezes the language backbone during training, shows improved instruction-following capabilities.
@@ 4.9 Discussion: Architecture and Data Contributions to Generalization
    LIBERO (Liu et al., 2023a): Evaluates the model’s capacity to fit training distributions and execute long-horizon, multi-stage manipulation, stressing policy learning and temporal credit assignment.
@@
    WidowX Generalization with BridgeData V2 (Walke et al., 2023): Five task groups selected from the WidowX 250 S Prepare Vegetables and Arrange Cups suites. We adopt the generalization protocols in Section˜4.7: Visual (changes in object appearance, background, lighting), Action (altered object positions, novel viewpoints, new instances), and Language (typos, rephrasings, fuzzy descriptions). For training, we randomly sample relevant demonstrations from BridgeData at four scales: 50, 500, 5k, and 50k.
@@
    -1 (fast): A fully autoregressive baseline using the Fast-tokenizer (Pertsch et al., 2025), trained on Bridge Data with interleaved vision–text–action pairs.
@@ 5 Related Work
Vision-Language-Action Models. Vision-language-action models (VLAs) have emerged as a promising approach for generalizable robot control by extending pre-trained vision-language models (VLMs) to action prediction. Autoregressive approaches (Brohan et al., 2023, Kim et al., 2024, Qu et al., 2025, Pertsch et al., 2025, Liu et al., 2024b) discretize continuous actions into tokens with linear bins, spatial grids or quantized DCT byte-pair encoding, naturally aligning with VLM training paradigms and exhibiting inherent advantages in knowledge transfer from web-scale pre-training. Despite these advantages, this discrete paradigm suffers from inference latency and limited resolution for high-frequency control tasks. Alternative works (Octo et al., 2024, Liu et al., 2024b, Black et al., 2024, Bjorck et al., 2025) introduce continuous action experts via diffusion or flow-matching (Ho et al., 2020, Lipman et al., 2022, Song et al., 2025), achieving fast inference and dexterous robotic control. Several efforts (Liu et al., 2025, Kim et al., 2025) attempt to combine discrete autoregression and continuous denoising, employing collaborative action ensemble or two-stage tuning procedure. However, these models suffer from the challenges of knowledge preservation and training efficiency, as the gradients from the action expert degrade the pre-trained VLM backbone (Driess et al., 2025, Black et al., 2025).
@@
Co-training strategies have emerged as a promising solution that integrates diverse data sources and enables generalization to new environments. $\pi_{0.5}$ (Black et al., 2025) employs a two-stage co-training approach: first pre-training in multimodal data sources with FAST tokenized actions, then tuning the action expert for low-level control. Building on this foundation, (Driess et al., 2025) formalizes a single-stage recipe insulating the VLM backbone during VLA training. Complementary works (Lin et al., 2025, Zhou et al., 2025) further explore reasoning and generalization through $\pi_{0}$-style co-training or purely autoregressive paradigms. Nevertheless, instruction following capabilities remain suboptimal, fundamentally due to architectural bottlenecks and the fact that current approaches are predominantly trained on separate image-text pairs or robotic episodes data, overlooking the rich temporal dynamics and causal relationships inherent in embodied interactions. Addressing these challenges necessitates better co-training strategies, improved cross-modal transfer mechanisms, and larger-scale interleaved datasets that comprehensively capture the full spectrum of robotic episodes.
@@
Unified Multimodal Models. Unifying understanding and generation (Deng et al., 2025, Chen et al., 2025b, Liao et al., 2025) has witnessed remarkable progress and emergent properties beyond individual specialized models. Pioneering unified multimodal models (Zhou et al., 2024, Team, 2024, Xie et al., 2024) integrate these capabilities through autoregressive or diffusion modeling in a single architecture, fusing both text autoregression and visual generation in the shared backbone. Recent advancements (Deng et al., 2025, Ma et al., 2025a, Chen et al., 2025b, Xie et al., 2025) introduce modality-specific paramaters or dual visual encoding to optimize training pipelines, facilitate joint token semantic and superior performance. Alternative studies (Chen et al., 2025a, Dong et al., 2023, Ge et al., 2024, Pan et al., 2025) explored unified models with frozen LLMs, assemble LMMs backbones and visual generative models via adapters or learnable queries. While assembling frozen LLMs struggles to demonstrate promising capabilities in robotic learning (Kim et al., 2024, Bjorck et al., 2025, Qu et al., 2025), unifying understanding and generation inspires architecture design in representative works (Black et al., 2024, 2025, Bjorck et al., 2025) and exhibits significant emergent open-world capabilities.
@@ 6 Conclusion and Future Work
In this article, we present EO-Robotics, a fully open training recipe to foster the research community to develop an advanced embodied foundation model. EO-Robotics is composed of EO -1 model, EO -Data1.5M dataset, and EO -Bench benchmark, where EO -Data1.5M is collected by a scalable pipeline for curating interleaved vision-text-action data to improve EO -1 model’s open-world generalization. EO -1 effectively synergizes auto-regressive decoding and flow matching denoising in single unified model that is co-trained with web data and curated interleaved embodied data EO -Data1.5M, enabling seamless perception, planning, reasoning, and acting in complex physical environments. Our evaluations across embodied reasoning and robot control experiments demonstrate that EO -1 consistently outperforms strong vision–language–action baselines, exhibiting superior multimodal reasoning and dexterous robot control capabilities in open-world generalization. Moreover, EO -Bench presents a new benchmark to evaluate embodied reasoning capabilities of embodied foundation models. In summary, EO-Robotics is fully open-sourced to the community and achieves a forward step for equipping general-purpose autonomous robots with human-like reasoning and acting abilities.
EO-Robotics has made a solid step toward developing general embodied AI, featuring with seamless embodied reasoning and action in the open world. While the results with EO -1 demonstrate promising robot control capabilities, there remains key aspects that future work can address. First, we aim to enhance EO -1’s reasoning and action ability to handle complex scenarios involving navigation, obstacle avoidance, failure detection and analysis, human intent recognition, and human-robot interaction/cooperation. This requires seamlessly integrate more general multimodal understanding, embodied reasoning, and action abilities into one system, leading to more easy-to-use robots in real life. Second, we plan to explore more efficient design of unified model architecture or asynchronous inference pipeline to enable human-like simultaneous reasoning and action. Finally, we will incorporate more data and robot embodiments (e.g., human data and humanoid robot) into training recipe to enable EO-Robotics to be a boarder foundation for general-purpose autonomous robots.
@@ D.3.1 Temporal Reasoning
For temporal reasoning, we focus on two key aspects: task planning and physical common sense. Task planning questions require understanding the sequential structure of subtasks, while physical common sense questions evaluate the model’s ability to reason about object dynamics and feasibility in the physical world. Figure˜28 and 29 provide examples of the prompts used to elicit these two categories of questions.
@@ D.3.2 Spatial Reasoning
For spatial reasoning, we construct four sub-tasks: trajectory prediction, object pointing, object referring, and multiview reasoning. The multiview data are generated using the same prompting strategy as object pointing, without a separate prompt design. Figure˜30, 31, and 32 show representative prompts used to construct these spatial reasoning tasks. Together, they ensure that the dataset captures fine-grained spatial relations necessary for embodied intelligence.
Figure 30: Trajectory prediction prompt.

### v5

@@ 4.3 Benchmark Evaluation Results
RT-1 (brohan2022rt) 85.7% 44.2% 73.0% 6.5% 52.4%
OpenVLA (kim2024openvla) 16.3% 46.2% 35.6% 0% 24.5%
TraceVLA (zheng2024tracevla) 28.0% 53.7% 57.0% 0% 34.7%
SpatialVLA (qu2025spatialvla) 86.0% 77.9% 57.4% 0% 55.3%
Magma (yang2025magma) 75.0% 53.0% 58.9% 8.3% 48.8%
$\pi_{0}$-FAST (pertsch2025fast) 75.3% 67.5% 42.9% 0 46.4%
RT-1-X (o2024open) 56.7% 31.7% 59.7% 40.7% 47.2%
RT-2-X (o2024open) 78.7% 77.9% 25.0% 7.4% 47.3%
Octo-Base (team2024octo) 17.0% 4.2% 22.7% 0 11.0%
RoboVLM li2023generalist 77.3% 61.7% 43.5% 24.1% 51.7%
$\pi_{0}$ (black2024pi_0) 97.9% 78.7% 62.25% 46.6% 71.4%
ThinkAct (huang2025thinkact) 92.0% 72.4% 50.0%
MolmoAct (lee2025molmoact) 77.7% 77.1% 60.0%
@@
RT-1 (brohan2022rt) 89.8% 50.0% 32.3% 2.6% 43.7%
OpenVLA (kim2024openvla) 54.5% 47.7% 17.7% 0.0% 30.0%
TraceVLA (zheng2024tracevla) 60.0% 56.4% 31.0% 0% 36.9%
SpatialVLA (qu2025spatialvla) 88.0% 72.7% 41.8% 6.3% 52.2%
Magma (yang2025magma) 68.6% 78.5% 59.0% 24.0% 57.5%
$\pi_{0}$-FAST (pertsch2025fast) 77.6% 68.2% 31.3% 0 44.3%
RT-1-X (o2024open) 49.0% 32.3% 29.4% 10.1% 30.2%
RT-2-X (o2024open) 82.3% 79.2% 35.3% 20.6% 54.4%
Octo-Base (team2024octo) 0.6% 3.1% 1.1% 0 1.2%
RoboVLM li2023generalist 75.6% 60.0% 10.6% 0 36.6%
$\pi_{0}$ (black2024pi_0) 90.1% 80.7% 27.6% 20.5% 54.7%
ThinkAct (huang2025thinkact) 84.0% 63.8% 47.6%
@@
Table 4: Performance comparison with state-of-the-art policies on SimplerEnv Benchmark (li24simpler).
Robot Control. We evaluated manipulation performance on two simulation benchmarks: LIBERO (liu2023libero)and SimplerEnv (li24simpler). LIBERO spans four subsets, Spatial, Object, Goal, and Long, designed to test generalization across spatial layouts, object categories, goal semantics, and long-horizon planning. SimplerEnv includes Google-VM (Visual Matching), Google-VA (Visual Aggregation), and WindoWX setups with controlled variations in color, texture, lighting, and camera pose, assessing robustness under visual distribution shift. We compare against leading robot foundation models, including the RT series (brohan2022rt, o2024open), Octo (team2024octo), OpenVLA (kim2024openvla), the $\pi$-series (black2024pi_0, pertsch2025fast), and ThinkAct (huang2025thinkact), among others.
As shown in Table˜3 and Table˜4, our method achieves superior performance across various robotic manipulation tasks. On the LIBERO benchmark, it attains an average success rate of 98.2%, outperforming recent state-of-the-art models, including OpenVLA-OFT (kim2025fine), $\pi_{0}$ (black2024pi_0), and GR00t-N1 (bjorck2025gr00t) by 1.1%, 4.0%, and 4.3%, respectively, demonstrating robust generalization across spatial, semantic, and long-horizon tasks. On the SimplerEnv benchmarks, it surpasses $\pi_{0}$ (black2024pi_0) by 3.5%, 5.1%, and 8.3% in the WidowX, Google-VM, and Google-VA variants, respectively, achieving the highest success rates of 72.7%, 76.5%, and 63.0%. These results are obtained with lightweight fine-tuning on a modest mixed-modality dataset, underscoring the method’s data efficiency, dexterous control, and precise visuomotor grounding.
@@ 4.5 Mastering Diverse Manipulations on Multiple Embodiments
In this section, we demonstrate that EO -1 can perform a wide range of dexterous manipulation tasks across various robotic platforms, showcasing its robustness and adaptability. We evaluate its performance on both short-horizon and long-horizon tasks using different robots, including Franka Panda, WidowX 250 S, Agibot G-1, and Lerobot SO100. These tasks range from basic pick-and-place to complex articulated object manipulation and long-horizon tasks. We compare EO -1 against state-of-the-art baselines, including $\pi_{0}$-Fast, $\pi_{0}$, and GR00T-N1.5 (pertsch2025fast, black2024pi_0, bjorck2025gr00t), highlighting its superior performance across these varied tasks.
@@ 4.7 Emerging Open-world Embodied Generalization
Against strong baselines GR00T-N1.5 and $\pi_{0}$ (pertsch2025fast, black2024pi_0, bjorck2025gr00t), EO -1 achieves an average completion score of 87.0%, outperforming GR00T-N1.5 (75.0%) and $\pi_{0}$ (66.0%). Performance gains are consistent across embodiments and task types, with EO -1 maintaining high accuracy even under precise spatial references and unseen object–instruction combinations, demonstrating robustness to both language and visual distribution shifts. We observe that while $\pi_{0}$ executes tasks quickly and accurately, it struggles to follow instructions, often prioritizing objects closer to itself instead of adhering to the specified commands. This behavior may be due to the negative impact of pre-training on the entire backbone, which can lead to suboptimal decision-making in dynamic environments. In contrast, GR00T-N1.5, which freezes the language backbone during training, shows improved instruction-following capabilities.
@@ 4.9 Discussion: Architecture and Data Contributions to Generalization
    LIBERO (liu2023libero): Evaluates the model’s capacity to fit training distributions and execute long-horizon, multi-stage manipulation, stressing policy learning and temporal credit assignment.
@@
    WidowX Generalization with BridgeData V2 (walke2023bridgedata): Five task groups selected from the WidowX 250 S Prepare Vegetables and Arrange Cups suites. We adopt the generalization protocols in Section˜4.7: Visual (changes in object appearance, background, lighting), Action (altered object positions, novel viewpoints, new instances), and Language (typos, rephrasings, fuzzy descriptions). For training, we randomly sample relevant demonstrations from BridgeData at four scales: 50, 500, 5k, and 50k.
@@
    -1 (fast): A fully autoregressive baseline using the Fast-tokenizer (pertsch2025fast), trained on Bridge Data with interleaved vision–text–action pairs.
@@ 5 Related Work
Vision-Language-Action Models. Vision-language-action models (VLAs) have emerged as a promising approach for generalizable robot control by extending pre-trained vision-language models (VLMs) to action prediction. Autoregressive approaches (brohan2023rt, kim2024openvla, qu2025spatialvla, pertsch2025fast, liu2024rdt) discretize continuous actions into tokens with linear bins, spatial grids or quantized DCT byte-pair encoding, naturally aligning with VLM training paradigms and exhibiting inherent advantages in knowledge transfer from web-scale pre-training. Despite these advantages, this discrete paradigm suffers from inference latency and limited resolution for high-frequency control tasks. Alternative works (team2024octo, liu2024rdt, black2024pi_0, bjorck2025gr00t) introduce continuous action experts via diffusion or flow-matching (ho2020denoising, lipman2022flow, song2025hume), achieving fast inference and dexterous robotic control. Several efforts (liu2025hybridvla, kim2025fine) attempt to combine discrete autoregression and continuous denoising, employing collaborative action ensemble or two-stage tuning procedure. However, these models suffer from the challenges of knowledge preservation and training efficiency, as the gradients from the action expert degrade the pre-trained VLM backbone (driess2025knowledge, intelligence2025pi_0_5).
@@
Co-training strategies have emerged as a promising solution that integrates diverse data sources and enables generalization to new environments. $\pi_{0.5}$ (intelligence2025pi_0_5) employs a two-stage co-training approach: first pre-training in multimodal data sources with FAST tokenized actions, then tuning the action expert for low-level control. Building on this foundation, (driess2025knowledge) formalizes a single-stage recipe insulating the VLM backbone during VLA training. Complementary works (lin2025onetwovla, zhou2025chatvla) further explore reasoning and generalization through $\pi_{0}$-style co-training or purely autoregressive paradigms. Nevertheless, instruction following capabilities remain suboptimal, fundamentally due to architectural bottlenecks and the fact that current approaches are predominantly trained on separate image-text pairs or robotic episodes data, overlooking the rich temporal dynamics and causal relationships inherent in embodied interactions. Addressing these challenges necessitates better co-training strategies, improved cross-modal transfer mechanisms, and larger-scale interleaved datasets that comprehensively capture the full spectrum of robotic episodes.
@@
Unified Multimodal Models. Unifying understanding and generation (deng2025emerging, chen2025janus, liao2025mogao) has witnessed remarkable progress and emergent properties beyond individual specialized models. Pioneering unified multimodal models (zhou2024transfusion, team2024chameleon, xie2024show) integrate these capabilities through autoregressive or diffusion modeling in a single architecture, fusing both text autoregression and visual generation in the shared backbone. Recent advancements (deng2025emerging, ma2025unitok, chen2025janus, xie2025show) introduce modality-specific paramaters or dual visual encoding to optimize training pipelines, facilitate joint token semantic and superior performance. Alternative studies (chen2025blip3, dong2023dreamllm, ge2024seed, pan2025transfer) explored unified models with frozen LLMs, assemble LMMs backbones and visual generative models via adapters or learnable queries. While assembling frozen LLMs struggles to demonstrate promising capabilities in robotic learning (kim2024openvla, bjorck2025gr00t, qu2025spatialvla), unifying understanding and generation inspires architecture design in representative works (black2024pi_0, intelligence2025pi_0_5, bjorck2025gr00t) and exhibits significant emergent open-world capabilities.
@@ 6 Conclusion and Future Work
In this article, we present EO -1, a fully open training recipe to foster the research community to develop an advanced embodied foundation model. EO -1 is composed of EO -1 model, EO -Data1.5M dataset, and EO -Bench benchmark, where EO -Data1.5M is collected by a scalable pipeline for curating interleaved vision-text-action data to improve EO -1 model’s open-world generalization. EO -1 effectively synergizes auto-regressive decoding and flow matching denoising in a single unified model that is co-trained with web data and curated interleaved embodied data EO -Data1.5M, enabling seamless perception, planning, reasoning, and acting in complex physical environments. Our evaluations across embodied reasoning and robot control experiments demonstrate that EO -1 consistently outperforms strong vision–language–action baselines, exhibiting superior multimodal reasoning and dexterous robot control capabilities in open-world generalization. Moreover, EO -Bench presents a new benchmark to evaluate embodied reasoning capabilities of embodied foundation models. In summary, EO -1 is fully open-sourced to the community and achieves a forward step for equipping general-purpose autonomous robots with human-like reasoning and acting abilities.
EO
-1 has made a solid step toward developing general embodied AI, featuring with seamless embodied reasoning and action in the open world. While the results with EO -1 demonstrate promising robot control capabilities, there remains key aspects that future work can address. First, we aim to enhance EO -1’s reasoning and action ability to handle complex scenarios involving navigation, obstacle avoidance, failure detection and analysis, human intent recognition, and human-robot interaction/cooperation. This requires seamlessly integrate more general multimodal understanding, embodied reasoning, and action abilities into one system, leading to easier-to-use robots in real life. Second, we plan to explore a more efficient design of a unified model architecture or an asynchronous inference pipeline to enable human-like simultaneous reasoning and action. Finally, we will incorporate more data and robot embodiments (e.g., human data and humanoid robots) into the training recipe to enable EO -1 to be a broader foundation for general-purpose autonomous robots.
@@ D.3.1 Temporal Reasoning
For temporal reasoning, we focus on two key aspects: task planning and physical common sense. Task planning questions require understanding the sequential structure of subtasks, while physical common sense questions evaluate the model’s ability to reason about object dynamics and feasibility in the physical world. Figure˜28 and 30 provide examples of the prompts used to elicit these two categories of questions.
@@ D.3.2 Spatial Reasoning
For spatial reasoning, we construct four sub-tasks: trajectory prediction, object pointing, object referring, and multiview reasoning. The multiview data are generated using the same prompting strategy as object pointing, without a separate prompt design. Figure˜29, 31, and 32 show representative prompts used to construct these spatial reasoning tasks. Together, they ensure that the dataset captures fine-grained spatial relations necessary for embodied intelligence.